Compute the magnitude of the moment of the force P about point A


1
Expert's answer
2021-05-31T17:02:38-0400

Gives

Point A 

Magnitude of moment of force=P


The moment of a force about a point provides a measure of the tendency for rotation (sometimes called a torque). As shown,

d is the perpendicular distance from point A to the line of action of the force. by using the components of F as shown.

Moment of force =force"\times" perpendicular distance

"\tau=F\times d"

2-D moment of force

"\tau=\begin{vmatrix}
\hat i&\hat j&\hat k\\
 F_x & F_y &F_z\\
 d_x&d_y&d_z
\end{vmatrix}"


"\tau=\hat i (F_yd_z-F_zd_y)-\hat j(F_xd_z-F_zd_x)+\hat k(F_xd_y-F_yd_x)"